There are 5 defective light bulbs in a bin containing 60 light bulbs.How many ways can you choose at least 4 defective light bulbs at random

1
Expert's answer
2022-02-07T15:16:51-0500

4 defective and 2 non-defective bulbs could be selected in "C_5^4*C_{55}^2=7425\;ways."

5 defective and 1 non-defective bulbs could be selected in "C_5^5*C_{55}^1=55\;ways."

So the number of ways at least 4 defective bulbs can be selected when 6 bulbs are drawn randomly:

"N=7425+55=7480."